# OmniNWM: Omniscient Driving Navigation World Models

### v3

@@ Appendix C Semantic and Depth Data Curation
To construct semantic labels for the NuScenes dataset [6], we finetune a pretrained DDRNet model [27] on diverse driving scenes to enhance robustness. Training data includes the Cityscapes [12], Mapillary Vistas [61], Waymo Open [68], Woodscape [101], and BDD100k [102] datasets. Image-to-image translation techniques [35] are employed to synthesize night-scene imagery, further improving model generalization. The finetuned DDRNet generates semantic segmentation labels for the NuScenes dataset at 12 Hz.
@@ Appendix E Training Details
Stage 3: Variable-length and Resolution Training. Following [115], we introduce variability in sequence length (17 or 33 frames) and resolution ($224\times 400$ or $448\times 800$) to improve adaptability to diverse driving scenarios and computational constraints.
@@ Appendix F More Visualization Results
Trajectory Planning Evaluation. Figure 12 presents the average rewards for specific scenarios on the NuScenes validation set across VLA planning baselines, evaluated using our occupancy-grounded driving reward. The results indicate that our reward function effectively distinguishes between the planning trajectories generated by different VLAs—such as those produced when the ego vehicle navigates near drivable area boundaries of median strips, or oncoming trucks.

### v4

@@ Appendix C Semantic and Depth Data Curation
To construct semantic labels for the NuScenes dataset [6], we finetune a pretrained DDRNet model [27] on diverse driving scenes to enhance robustness. Training data includes Cityscapes [12], Mapillary Vistas [61], Waymo Open [68], Woodscape [101], and BDD100k [102] datasets. Image-to-image translation techniques [35] are employed to synthesize night-scene imagery, further improving model generalization. The finetuned DDRNet generates semantic segmentation labels for the NuScenes dataset at 12 Hz.
@@ Appendix E Training Details
Stage 3: Variable-length and Resolution Training (3K iterations). Following [115], we introduce variability in sequence length (17 or 33 frames) and resolution ($224\times 400$ or $448\times 800$) to improve adaptability to diverse driving scenarios and computational constraints.
@@ Appendix F More Visualization Results
Trajectory Planning Evaluation. Figure 12 presents the average rewards for specific scenarios on the NuScenes validation set across VLA planning baselines, evaluated using our occupancy-grounded driving reward. The results indicate that our reward function effectively distinguishes between the planning trajectories generated by different VLAs when the ego vehicle navigates near drivable area boundaries of median strips, or oncoming trucks.